# Learning Encoding-Decoding Direction Pairs to Unveil Concepts of Influence in Deep Vision Networks

### v1

@@ A.3 Direction Learning Process
We learn the directions of the proposed method in a four-step process: a) we first learn the parameters ${\bm{w}}_{i},b_{i}$ following Doumanoglou et al. (2024), replacing the CNN Classifier Loss with the proposed $\mathcal{L}^{uur}$; b) we then continue optimizing ${\bm{w}}_{i},b_{i}$, removing the orthogonality and standardization constraints while keeping $\mathcal{L}^{uur}$ and incorporating the additional losses from Section 4.2; c) next, we learn the signal vectors using the filters of the learned classifiers as regressors in (1) to initialize $\{\hat{{\bm{S}}}\}$; and d) finally, we jointly optimize $\hat{{\bm{s}}}_{i}$, ${\bm{w}}_{i},b_{i}$, using all previous losses, replacing $\mathcal{L}^{uur}$ with $\mathcal{L}^{cur}$ and adding $\mathcal{L}^{fso}$ from Sections 4.3 and 4.4.
A.4 Details for the Experiment on Synthetic Data
@@
We train the network using cross-entropy loss and the Adam (Kingma, 2014) optimizer, with learning rate $0.0005$ and batch size $1024$ for $4000$ epochs. In principle, we follow the process defined in Section A.3, but due to the simplicity of the example, during (a) we don’t use $\mathcal{L}^{uur}$, we omit step (b) and proceed directly from (a) to (c). In both steps (a) and (d) we use the Augmented Lagrangian Loss, the Adam optimizer, and the Cosine Annealing learning rate scheduler (Loshchilov & Hutter, 2016). In step (a), we omit the constraints of the Augmented Lagrangian loss that are not applicable. The Augmented Lagrangian formulation greatly stabilizes learning and avoids local optima. For step (a) we solve the constrained optimization problem of minimizing $\mathcal{L}^{s}$ with $\tau^{ma}=0.8$, $\tau^{mm}=5$ and $\tau^{ic}=0.0$. For step (d) we minimize $\lambda^{fs}\mathcal{L}^{fs}+\lambda^{cur}\mathcal{L}^{cur}$ with $\tau^{ma}=0.5$, $\tau^{mm}=15$, $\tau^{ic}=0.0$, $\tau^{eac}=0.0$, $\tau^{fso}=0.01$. The learning rate we use for step (a) is $0.00025$ and for step (b) is $0.0005$. The number of learning epochs is set to $10000$ and $20000$, for each step respectively. For the loss weights we use $\lambda^{fs}=2.6$ and $\lambda^{ur}=0.25$. The sharpening factor $\gamma$ of $\mathcal{L}^{ic}$ is set to $\gamma=2.0$, the $\tau$ hyperparameter of the same loss is set to $\tau=1.0$, the $\nu$ sharpening factor of $\mathcal{R}_{SW}$ to $\nu=2.0$, the $\rho$ of $\mathcal{L}^{eac}$ is set to $\rho=5/3$ and the $\mu$ sharpening factor of $\mathcal{L}^{fs}$ is set to $2.0$.
@@
A.5 Extracting Signal Values with the Filters of Concept Detectors
@@
A.6 Details on the Faithfulness Assessment of Encoding-Decoding Direction Pairs
@@
Due to the considerable computational load required to conduct this experiment, we make some simplifications. First, we consider a subset of the learned direction pairs by randomly choosing $100$. Candidate pairs are the ones with a concept detector that is considered interpretable by Network Dissection (i.e. the detector exhibits an IoU performance greater than $0.04$ when detecting the concept). Second, to construct the image set that contains the concept, we consider $10$ unique images within the set of the most confident (patch-level) predictions of the concept detector. For each image, we run the deep dream optimization for $K=40$ iterations, and we record the patch embeddings (feature activations) containing the concept with an interval of $5$ steps. For dreaming, we don’t use zooming, but we do use the robustness transforms originally proposed in Olah et al. (2017). Regarding the line fitting process, we consider a parametric line as ${\bm{c}}+\lambda{\bm{d}}$, $\lambda\in{\mathbb{R}}$, with ${\bm{c}}\in{\mathbb{R}}^{D}$ denoting the line’s origin and ${\bm{d}}\in{\mathbb{R}}^{D}$ its direction (the dreaming direction). At the end of the dreaming optimization loop and after recording the feature activations of the concept, we minimize the average distance from each feature to the line. Finally, we make sure that the dreaming direction points towards the direction of feature evolvement across iterations by considering making a sign flip, i.e. multiplying $\bm{d}$} with 1-1.
A.7 Details for the Experiments on Deep Image Classifiers
@@ II Step Epochs LR
Learning Details In Table 15 we provide the main hyper-parameters to train the direction pairs for each step. The learning rate that we provide (LR) is considered as the reference learning rate for a batch size of 4096. In practice, we scale both the learning rate and the batch size based on the available GPU memory and the number of GPUs. We use the Adam Kingma (2014) optimizer and the cosine annealing learning rate scheduler (Loshchilov & Hutter, 2016). While for step (a) we follow the hyper-parameter setup of Doumanoglou et al. (2024), for steps (b) and (d) we make different choices. For $\mathcal{L}^{ic}$ and $\mathcal{L}^{eac}$ we set their hyper-parameters $\tau$ and $\rho$ as follows: For ResNet18 and VGG16, $\tau$ and $\rho$ are set such that the minimum cluster size is equal to 400 and the maximum 50000. For EfficientNet and Inception-v3, the same numbers are 520 and 65000, while for ResNet50 the minimum cluster size is set to 920 and the maximum to 75000. These are all set based on the statistics of the concept datasets. The $\mu$ sharpening factor of $\mathcal{L}^{fs}$ and the $\nu$ sharpening factor of $\mathcal{R}_{SW}$ are set to $2.0$.
@@
When using $\mathcal{L}^{uur}$ and $\mathcal{L}^{cur}$, we observed better results when manipulating features with a stochastic magnitude in the direction $\bm{dx_{{\bm{p}}}}$, i.e. shifting representations as ${\bm{x}}_{{\bm{p}}}^{\prime}={\bm{x}}_{{\bm{p}}}-\kappa\bm{dx_{{\bm{p}}}}$ with $\kappa$ a random number in $[0.1,0.5]$. In practice, we separate filter directions from their magnitude $1/M_{i}$ and learn them independently as suggested in Doumanoglou et al. (2024). For enforcing ${\bm{w}}_{i}||_{2}=1$ (i.e. unit norm filter vectors) we use parametrization on the unit hyper-sphere.
@@
Details on Datasets used to Compute Concept Sensitivity with RCAV To conduct concept sensitivity testing with RCAV we need access to a labeled dataset coming from the domain of the model. In practice, it is common to use the validation split of the dataset used to train the model. To mitigate the required computation time, we limited the size of the validation datasets as follows: For ImageNet, we used the validation split of ImageNet-S-300 Gao et al. (2022), and for Moments-in-Time, we considered Moments-In-Time Mini. Since Moments-In-Time is a video dataset, but the model that we studied works in the image domain, we constructed an image dataset by sampling 3 frames equally apart from each video (i.e. the first, the mid, and the last frame of the video). For Places365, we didn’t make a size reduction as the validation split of the dataset was manageable with our resources.
@@
Details on RCAV’s Statistical Significance Test In all experiments, we set RCAV’s perturbation hyper-parameter, to $\alpha=5$. For direction significance testing, we use RCAV’s label permutation test. To construct random noise signal vectors, we (a) construct a dataset of feature-(binary) label pairs based on the decision rule of each one of the concept detectors. To deal with great class imbalance, we construct a pool of negative samples that is at most 20 times more than the positive ones; (b) we construct $N$ noisy versions of that dataset by label permutation; (c) we learn a noise-classifier to distinguish features based on the permuted labels, and (d) we concurrently estimate a noise-signal vector using (1) and the subsampling process described in Section 4.3. To learn each one of the noise signal vectors, and before permuting the labels, we construct a balanced dataset of at most 5000 samples, picked randomly from the pool. We train the noise classifiers using Adam for 100 epochs and a learning rate $0.01$. By using noise signal vectors as RCAV’s noisy directions, and with the number of those vectors per classifier set to $N=100$, we subsequently calculate RCAV’s p-values. We apply Bonferroni correction to all p-values by dividing the significance threshold $0.05$ by the number of concept detectors $I$ and the number of model classes.
@@
Learning Details for NMF When learning directions with NMF, we used the implementation found in Yu (2020). To make computations manageable, we reduce the spatial dimensionality of image representations by applying Adaptive Average Pooling to a (2,2) resolution, in a similar way as it was done in Fel et al. (2023c). In all cases, we learn the NMF decomposition by running the algorithm for $100000$ iterations.
@@
Details for Comparing with the Supervised Approach When learning directions with IBD, for each concept, a dataset is assembled that includes up to 20 times more negative samples than positive ones to address the significant imbalance. Additionally, we use hard negative mining as originally suggested in Zhou et al. (2018).
@@
Figure 23: Interpretability Comparison. Histogram of differences in binary metrics: Precision, Recall, F1Score between the Linear-OR set of concept detectors learned with EDDP-C, $I=512$, and classifiers learned in a supervised way (IBD Zhou et al. (2018)). The network here is ResNet18 trained on Places365.
@@
Figure 23 plots a histogram of classification metric differences between the Linear-OR set of classifiers and the classifiers learned in a supervised way. The differences are based on the concept labels, effectively taking the difference of metrics that regard two classifiers (the first from the Linear-OR set and the second from Zhou et al. (2018)) with the same concept name.
@@
Figures 24, 25, 26 depict concrete binary classification metrics for some of the concept detectors in the Linear-OR set of classifiers, comparing them with concept detectors learned with supervision.
@@
Figure 24: Interpretability Comparison. Exact Precision/Recall/F1Scores for specific concepts in Broden: comparison between the Linear-OR set of classifiers learned with EDDP-C, $I=512$, and classifiers learned in a supervised way (IBD Zhou et al. (2018)). The network here is ResNet18 trained on Places365.
@@
Figure 25: Interpretability Comparison. Exact Precision/Recall/F1Scores for specific concepts in Broden: comparison between the Linear-OR set of classifiers learned with EDDP-C, $I=512$, and classifiers learned in a supervised way (IBD Zhou et al. (2018)). The network here is ResNet18 trained on Places365.
@@
Figure 26: Interpretability Comparison. Exact Precision/Recall/F1Scores for specific concepts in Broden: comparison between the Linear-OR set of classifiers learned with EDDP-C, $I=512$, and classifiers learned in a supervised way (IBD Zhou et al. (2018)). The network here is ResNet18 trained on Places365.
@@
A.8 Hyper-Parameter Study with Respect to Target Separation Margin

### v2

@@ A.7 Direction Learning Process
We learn the directions of the proposed method in a four-step process: a) we first learn the parameters ${\bm{w}}_{i},b_{i}$ following CBE, replacing the CNN Classifier Loss with the proposed $\mathcal{L}^{uur}$; b) we then continue optimizing ${\bm{w}}_{i},b_{i}$, removing the orthogonality and standardization constraints while keeping $\mathcal{L}^{uur}$ and incorporating the additional losses from Section 4.2; c) next, we learn the signal vectors using the filters of the learned classifiers as signal value extractors in (1) to initialize $\{\hat{{\bm{S}}}\}$; and d) finally, we jointly optimize $\hat{{\bm{s}}}_{i}$, ${\bm{w}}_{i},b_{i}$, using all previous losses, replacing $\mathcal{L}^{uur}$ with $\mathcal{L}^{cur}$ and adding $\mathcal{L}^{fso}$ from Sections 4.3 and 4.4.
A.8 Details for the Experiment on Synthetic Data
@@
We train the network using cross-entropy loss and the Adam (adam) optimizer, with learning rate $0.0005$ and batch size $1024$ for $4000$ epochs. In principle, we follow the process defined in Section A.7, but due to the simplicity of the example, during (a) we don’t use $\mathcal{L}^{uur}$, we omit step (b) and proceed directly from (a) to (c). In both steps (a) and (d) we use the Augmented Lagrangian Loss, the Adam optimizer, and the Cosine Annealing learning rate scheduler (cosine-annealing). In step (a), we omit the constraints of the Augmented Lagrangian loss that are not applicable. The Augmented Lagrangian formulation greatly stabilizes learning and avoids local optima. For step (a) we solve the constrained optimization problem of minimizing $\mathcal{L}^{s}$ with $\tau^{ma}=0.8$, $\tau^{mm}=5$ and $\tau^{ic}=0.0$. For step (d) we minimize $\lambda^{fs}\mathcal{L}^{fs}+\lambda^{cur}\mathcal{L}^{cur}$ with $\tau^{ma}=0.5$, $\tau^{mm}=15$, $\tau^{ic}=0.0$, $\tau^{eac}=0.0$, $\tau^{fso}=0.01$. The learning rate we use for step (a) is $0.00025$ and for step (b) is $0.0005$. The number of learning epochs is set to $10000$ and $20000$, for each step respectively. For the loss weights we use $\lambda^{fs}=2.6$ and $\lambda^{ur}=0.25$. The sharpening factor $\gamma$ of $\mathcal{L}^{ic}$ is set to $\gamma=2.0$, the $\tau$ hyperparameter of the same loss is set to $\tau=1.0$, the $\nu$ sharpening factor of $\mathcal{R}_{SW}$ to $\nu=2.0$, the $\rho$ of $\mathcal{L}^{eac}$ is set to $\rho=5/3$ and the $\mu$ sharpening factor of $\mathcal{L}^{fs}$ is set to $2.0$.
@@
A.9 Extracting Signal Values with the Filters of Concept Detectors
@@
A.10 Details on the Faithfulness Assessment of Encoding-Decoding Direction Pairs
@@
Due to the considerable computational load required to conduct this experiment, we make some simplifications. First, we consider a subset of the learned direction pairs by randomly choosing $100$. Candidate pairs are the ones with a concept detector that is considered interpretable by Network Dissection (i.e. the detector exhibits an IoU performance greater than $0.04$ when detecting the concept). Second, to construct the image set that contains the concept, we consider $10$ unique images within the set of the most confident (patch-level) predictions of the concept detector. For each image, we run the deep dream optimization for $K=40$ iterations, and we record the patch embeddings (feature activations) containing the concept with an interval of $5$ steps. For dreaming, we don’t use zooming, but we do use the robustness transforms originally proposed in feature-viz-olah. Regarding the line fitting process, we consider a parametric line as ${\bm{c}}+\lambda{\bm{d}}$, $\lambda\in{\mathbb{R}}$, with ${\bm{c}}\in{\mathbb{R}}^{D}$ denoting the line’s origin and ${\bm{d}}\in{\mathbb{R}}^{D}$ its direction (the dreaming direction). At the end of the dreaming optimization loop and after recording the feature activations of the concept, we minimize the average distance from each feature to the line. Finally, we make sure that the dreaming direction points towards the direction of feature evolvement across iterations by considering making a sign flip, i.e. multiplying $\bm{d}$} with 1-1.
A.11 Additional Experiments on the Faithfulness Assessment of the Encoding-Decoding Direction Pairs
Figures 23 and 24 complement the experiments presented in Figures 11 and 12 for the remaining combinations regarding $\mathcal{L}^{fso}$. Even though less prominent in ResNet18, together with Section 7.2 the experiments sufficiently support that $\mathcal{L}^{fso}$ pushes the distribution of cosine similarities between dreaming directions and signal vectors to become more bell-shaped. Overall, the experiments in this section, similar to the experiments presented in Section 7.2, indicate that in approximately 90% of the cases, these cosine similarities exceed 0.7. In a less rigorous but more qualitative sense, we found that when learning EDDP with $\mathcal{L}^{fso}$, this loss converges to values close to $\approx 0.02$, whereas measuring the loss value for a direction set learned without $\mathcal{L}^{fso}$ often leads to a value close to $\approx 0.1$, which is still not that far from indicating orthogonality between filters and signal vectors.
A.12 Details for the Experiments on Deep Image Classifiers
Details on Forming Concept Detectors While our approach explicitly learns binary classifiers $\{{\bm{w}}_{i},b_{i}\}$ as concept detectors, an estimation of an explicit classification threshold $b$ is missing when clustering with PCA, NMF, and the Natural basis. To form concept detectors in those cases, we follow NetworkDissection, and for each direction, we choose $b$ to be equal to the top $k$-quantile among the projected activations sourced from the concept dataset. For the Natural latent space basis and NMF, we use $k=0.005$, as suggested in NetworkDissection, while for PCA, we use $k=0.2$ as suggested in PCA.
Differences in the Evaluation Protocol compared to PCA-based and NMF-based State-of-the-Art and Limitations Previous methods (PCA; SVD; NMFCAV; CRAFT) based on PCA or NMF operated under a class-conditioned protocol, where the analysis and identified concepts are intrinsically linked to a specific class. On the contrary, in this work, our method is applied in a class-agnostic protocol, analyzing the latent space without prior class information. This fundamental difference in methodology renders a direct comparison with the exact implementations of these previous approaches (PCA; SVD; NMFCAV; CRAFT) unfair or infeasible. Specifically, as the size of the concept space $I$ approaches the size of the embedding space $D$, NMF trivially converges to the natural feature space basis. Moreover, in a class agnostic setting, the gradient re-ranking approach of (PCA; SVD) is not feasible, as gradients are tied to a specific class, requiring further adaptation of the original algorithm. However, PCA and NMF are still applicable in a class-agnostic protocol as typical matrix decomposition methods.
In this work, we implement both PCA and NMF baselines as matrix decomposition methods on layer feature activations in a class-agnostic fashion. We emphasize that, for NMF, we do not consider the recursivity part of CRAFT (as we consider concept factorization to be outside the scope of this work), and for PCA, we do not perform the gradient re-ranking that was suggested in PCA.
Given the previous limitations, both baselines could be less effective in the class-agnostic setting compared to the class-conditioned state-of-the-art. The challenges they face under the generalized, class-agnostic protocol highlight the necessity for more effective and novel approaches to identifying concept directions in this setting, and the proposed method is such an attempt.
Details on EDDP-U and EDDP-C We emphasize that EDDP-U does not include the use of $\mathcal{L}^{fso}$ to learn the directions. The main purpose of $\mathcal{L}^{fso}$ is to contribute to a faithful estimation of the encoding directions. While this was critical for the experiment on synthetic data, in real-world experiments, we see that the value of this loss is sufficiently low even when not directly minimizing it, possibly due to working in a high-dimensional embedding space where two random vectors are approximately orthogonal. A detailed discussion regarding $\mathcal{L}^{fso}$ is provided in Sections A.11 and A.14 with the conclusions presented in the main body of the article being mostly aligned with the conclusions that stem from detailed ablation studies.
Details on the Influence Evaluation Metric As we already mentioned in Section 3.5, RCAV originally used CAVs to assess the network’s concept sensitivity. Here we follow the proposition of PCAV and use the learned encoding directions in their place. For sensitivity scores, we spatially replicate the learned encoding directions across all spatial locations in the image representation. In studies without $\mathcal{L}^{cur}$, we report influence metrics for signal vectors that were estimated post learning the directions with the subsample strategy discussed in Section 4.3.
@@ II Step Epochs LR
Learning Details In Table 16 we provide the main hyper-parameters to train the direction pairs for each step. The learning rate that we provide (LR) is considered as the reference learning rate for a batch size of 4096. In practice, we scale both the learning rate and the batch size based on the available GPU memory and the number of GPUs. We use the Adam adam optimizer and the cosine annealing learning rate scheduler (cosine-annealing). While for step (a) we follow the hyper-parameter setup of CBE, for steps (b) and (d) we make different choices. For $\mathcal{L}^{ic}$ and $\mathcal{L}^{eac}$ we set their hyper-parameters $\tau$ and $\rho$ as follows: For ResNet18 and VGG16, $\tau$ and $\rho$ are set such that the minimum cluster size is equal to 400 and the maximum 50000. For EfficientNet and Inception-v3, the same numbers are 520 and 65000, while for ResNet50 the minimum cluster size is set to 920 and the maximum to 75000. These are all set based on the statistics of the concept datasets. The $\mu$ sharpening factor of $\mathcal{L}^{fs}$ and the $\nu$ sharpening factor of $\mathcal{R}_{SW}$ are set to $2.0$.
@@
When using $\mathcal{L}^{uur}$ and $\mathcal{L}^{cur}$, we observed better results when manipulating features with a stochastic magnitude in the direction $\bm{dx_{{\bm{p}}}}$, i.e. shifting representations as ${\bm{x}}_{{\bm{p}}}^{\prime}={\bm{x}}_{{\bm{p}}}-\kappa\bm{dx_{{\bm{p}}}}$ with $\kappa$ a random number in $[0.1,0.5]$. In practice, we separate filter directions from their magnitude $1/M_{i}$ and learn them independently as suggested in CBE. For enforcing ${\bm{w}}_{i}||_{2}=1$ (i.e. unit norm filter vectors) we use parametrization on the unit hyper-sphere.
@@
Details on Datasets used to Compute Concept Sensitivity with RCAV To conduct concept sensitivity testing with RCAV we need access to a labeled dataset coming from the domain of the model. In practice, it is common to use the validation split of the dataset used to train the model. To mitigate the required computation time, we limited the size of the validation datasets as follows: For ImageNet, we used the validation split of ImageNet-S-300 imagenet-s, and for Moments-in-Time, we considered Moments-In-Time Mini. Since Moments-In-Time is a video dataset, but the model that we studied works in the image domain, we constructed an image dataset by sampling 3 frames equally apart from each video (i.e. the first, the mid, and the last frame of the video). For Places365, we didn’t make a size reduction as the validation split of the dataset was manageable with our resources.
@@
Details on RCAV’s Statistical Significance Test In all experiments, we set RCAV’s perturbation hyper-parameter, to $\alpha=5$. For direction significance testing, we use RCAV’s label permutation test. To construct random noise signal vectors, we (a) construct a dataset of feature-(binary) label pairs based on the decision rule of each one of the concept detectors. To deal with great class imbalance, we construct a pool of negative samples that is at most 20 times more than the positive ones; (b) we construct $N$ noisy versions of that dataset by label permutation; (c) we learn a noise-classifier to distinguish features based on the permuted labels, and (d) we concurrently estimate a noise-signal vector using (1) and the subsampling process described in Section 4.3. To learn each one of the noise signal vectors, and before permuting the labels, we construct a balanced dataset of at most 5000 samples, picked randomly from the pool. We train the noise classifiers using Adam for 100 epochs and a learning rate $0.01$. By using noise signal vectors as RCAV’s noisy directions, and with the number of those vectors per classifier set to $N=100$, we subsequently calculate RCAV’s p-values. We apply Bonferroni correction to all p-values by dividing the significance threshold $p=0.05$ by the number of concept detectors $I$ and the number of model classes.
@@
Learning Details for NMF When learning directions with NMF, we used the implementation found in torch-nmf. To make computations manageable, we reduce the spatial dimensionality of image representations by applying Adaptive Average Pooling to a (2,2) resolution, in a similar way as it was done in CRAFT. In all cases, we learn the NMF decomposition by running the algorithm for $100000$ iterations, with $\text{beta}=2$ and a tolerance of 1e61e-6.
@@
Details for Comparing with the Supervised Approach When learning directions with IBD, for each concept, a dataset is assembled that includes up to 20 times more negative samples than positive ones to address the significant imbalance. Additionally, we use hard negative mining as originally suggested in IBD.
@@
Figure 25: Interpretability Comparison. Histogram of differences in binary metrics: Precision, Recall, F1Score between the Linear-OR set of concept detectors learned with EDDP-C, $I=512$, and classifiers learned in a supervised way (IBD IBD). The network here is ResNet18 trained on Places365.
@@
Figure 25 plots a histogram of classification metric differences between the Linear-OR set of classifiers and the classifiers learned in a supervised way. The differences are based on the concept labels, effectively taking the difference of metrics that regard two classifiers (the first from the Linear-OR set and the second from IBD) with the same concept name.
@@
Figures 26, 27, 28 depict concrete binary classification metrics for some of the concept detectors in the Linear-OR set of classifiers, comparing them with concept detectors learned with supervision.
@@
Figure 26: Interpretability Comparison. Exact Precision/Recall/F1Scores for specific concepts in Broden: comparison between the Linear-OR set of classifiers learned with EDDP-C, $I=512$, and classifiers learned in a supervised way (IBD IBD). The network here is ResNet18 trained on Places365.
@@
Figure 27: Interpretability Comparison. Exact Precision/Recall/F1Scores for specific concepts in Broden: comparison between the Linear-OR set of classifiers learned with EDDP-C, $I=512$, and classifiers learned in a supervised way (IBD IBD). The network here is ResNet18 trained on Places365.
@@
Figure 28: Interpretability Comparison. Exact Precision/Recall/F1Scores for specific concepts in Broden: comparison between the Linear-OR set of classifiers learned with EDDP-C, $I=512$, and classifiers learned in a supervised way (IBD IBD). The network here is ResNet18 trained on Places365.
@@
A.13 Hyper-Parameter Study with Respect to Target Separation Margin